how can I solve this solve each system of equations by substitution method
-3x+7y=14
2x-y=-13

Answer by lenny460(1073)

how can I solve this solve each system of equations by substitution method
-3x+7y=14
2x-y=-13

-3x + 7y = 14 (Equation 1)
2x - y = -13 (Equation 2)
2x - y = -13
Subtract 2x from each side of the equation.
2x - y - 2x = -13 - 2x
-y = -13 - 2x
Divide each item by -1
-y/-1 = -13/-1 -2x/-1
y = 13 + 2x or
y = 2x + 13 (Equation 3)
Substitute Equation 3 into Equation 1
-3x + 7y = 14
-3x + 7(2x + 13) = 14
-3x + 14x + 91 = 14
-3x + 14x = 11x
11x + 91 = 14
11x + 91 - 91 = 14 - 91
11x = -77
x = -7
Substitute x = -7 into Equation 2
2x - y = -13
2(-7) - y = -13
-14 - y = -13
Add 14 to each side of the equation.
-14 - y + 14 = -13 + 14
-y = 1
Divide each side by -1
y = -1

Answer:
x = -7 and
y = -1



Lennox Obuong
Algebra Tutor